OpenStudy (anonymous):

solve the eqn: cos(x+60)=2sinx)

OpenStudy (phi):

you could use cos(a+b)= cos(a)cos(b)- sin(a)sin(b) as a first step.

OpenStudy (anonymous):

use the double angle theroem

OpenStudy (anonymous):

cos60=1/2 and sin60=\[\cos(a+b)= \cos(a)\cos(b)- \sin(a)\sin(b) \] \[\cos(x)\cos(60)- \sin(x)\sin(60)=2\sin(x)\] \[\cos(60)= \frac{ 1 }{ 2} and \sin(60)=\frac{ \sqrt{3} }{2}\] \[\therefore\]\[\frac{ \cos(x) }{ 2 }-\frac{ \sqrt3(\sin(x)) }{ 2 }=2\sin(x)\]

OpenStudy (phi):

Obumpa, have you read the code of conduct? It would be better to let the asker make some progress on their own

OpenStudy (anonymous):

Oh sorry. You're right